OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

OpenStudy (hoblos):

general form y-b = m(x-a) where m is slope and (a,b) is a point belonging to the line

OpenStudy (hoblos):

can you do it now ?

OpenStudy (anonymous):

I'm not sure how to do this..

OpenStudy (hoblos):

try it y-(-1) = 6(x-5) so y+1=6(x-5)